OpenStudy (anonymous):

Dorothy is 14 years younger then Rita. Ten years ago, rita was 3 times as old as dorothy was then. how old is each now? step by step plz :)

OpenStudy (thurscon):

Dorothy = Rita - 14 (1) 3(Dorothy-10) = Rita-10 (2) Subs 1 into 2 3(Rita - 14 - 10) = Rita - 10 Rita - 24 = 1/3(Rita-10) Rita(1-1/3) = 24-10/3 Rita(2/3) = 62/3 Rita = 31 Dorothy = Rita - 14 Dorothy = 31-14 = 17 Rita is 31 years old and Dorothy is 17. Ten years ago, Dorothy was 7 and Rita was 21.